Example 7.2 Using method of joints find all forces in the truss below. ORA + 20k 1 T 120- 120- 121- 120- 121- 122-
Answer #1

Selution YE 12 124 12 ऋप At 1aint F924S6 FAG --44-22K Yain abou ea FAn=40k Aointe Fec -fhe FexKFen At 1oint Fan Afy E9.48 Put in abous ea 8-405 CsS Fn1=34K Teint C FeP 2 D Ce26 56fec Tos 26 + Pin 95 Fco =-38K Toint सु ु(-7.2)) Cs631+ as t39-a =-30k at EK 30k at@ 2.6 fatbn abo eg So.1K PEF fFEL-CK EL8-485 20 fik tFct feL Cr16K Taint € FDE 26-SE 45 FEF + FEf hin26 5 e in 26 FEL FEK t6 Fa hn FoE Co2 6t fef C 26- LFEK -K उ fER 44.3K Toint (K $6-31 f